OpenStudy (anonymous):

How far from the top must you cut a conical tent in order to cut the cloth in half?

OpenStudy (jdoe0001):

since there's no picture, I'd say half the height of the tent

OpenStudy (anonymous):

It can't be the half of the tent, though.. A cone when opened becomes a sector so there's probably another formula.. I just can't figure it out.

OpenStudy (bahrom7893):

 what's the surface area of the cone?

OpenStudy (anonymous):

It wasn't given.

OpenStudy (bahrom7893):

Well why don't you look it up?

OpenStudy (anonymous):

I tried to, but I couldn't understand the explanation. There also weren't much sources.

OpenStudy (jdoe0001):

do you know how to take a screenshot of the material?

OpenStudy (bahrom7893):

http://bit.ly/13delse

OpenStudy (anonymous):

Only the question was given.. there were no illustrations or what ever.

OpenStudy (jdoe0001):

hmmm, I notice the original posting was edited :P not it has "cloth" in it

OpenStudy (jdoe0001):

now rather

OpenStudy (bahrom7893):

@wlbadrm can you just go on google and search for the surface area of a cone?

OpenStudy (anonymous):

yeap, it was. I made a mistake typing the question. :)

OpenStudy (anonymous):

@bahrom7893 1/2(circumference of base)(slant height)

OpenStudy (jdoe0001):

total surface area of a cone = \(\bf \Large \pi r \sqrt{r^2+h^2} + \pi r^2\)

OpenStudy (bahrom7893):

we don't need the total surface area unless the cloth covers the bottom as well.

OpenStudy (jdoe0001):

\(\bf \sqrt{r^2+h^2}\) being the so-called slant height, using pythgorean theorem

OpenStudy (jdoe0001):

that's true,.... only lateral area , then just \(\bf \Large \pi r \sqrt{r^2+h^2}\)

OpenStudy (jdoe0001):

so I guess half of that Area for the tent, will be THAT/2 since no picture or values are provided, thus so it's

OpenStudy (bahrom7893):

So: \[A=\pi*r\sqrt{r^2+h^2}\]you want the area to be split into half. What should the new h be?